OpenStudy (mikezack123):

Admission to the movies costs $8 for adults and $4 for kids. The total bill for your entire extended family was $72. This scenario can be modeled by the equation 8a + 4k = 72. Which is not a possible combination of adults and kids? 3 adults and 12 kids 5 adults and 6 kids 7 adults and 4 kids 8 adults and 2 kids

OpenStudy (johnweldon1993):

Well you try them each out...for example Choice A) 3 adults and 12 kids....plug in 3 for 'a' and 12 for 'k' 8a + 4k = 72 8(3) + 4(12) = 72 24 + 48 = 72 72 = 72 So this is not your answer Try the others just like that...

OpenStudy (mikezack123):

i got C @johnweldon1993 right?

OpenStudy (johnweldon1993):

Sorry for the late response...no C is incorrect as well 7 adults and 4 kids so 8a + 4k = 72 8(7) + 4(4) = 72 56 + 16 = 72 72 = 72 So no C is not your answer

OpenStudy (johnweldon1993):

Lets look at B 5 adults and 6 kids 8a + 4k = 72 8(5) + 4(6) = 72 40 + 24 = 72 64 = 72 Here is your answer